Logbooks general detector daq Analysis Physics 12 GeV HallA OldDVCS2 DVCS2Ana Parity
runliste issues procurement survey minutes web compton2 NewIssues NewProcurement
  Minutes  Not logged in ELOG logo
Message ID: 6     Entry time: 06/23/04, 18:06
Author: Catherine Ferdi 
Subject: DVCS meeting minutes 
General Discussion:

- Tagger soon received in few days and guys from Grenoble coming next week
- Need organisation and device (oscilloscope, computer, coda...) in the EEL
for working (Tagger people, Proton array people and Calorimeter people)
- Dismount Calorimeter before moving it in the Hall A and mounting it in the
Hall A (It takes one week, no modification in the schedule)
- Bob Michaels in vacation after Happex and until 23 or 25/08 (Bodo could
replace him)
- 50 % beam time in addition ( all calculations done with 70% polarisation
and now for Happex 86% polarisation). Don't tell this to Kees and Larry...
- Need meeting with accelerator's guys after Happex (Sirish has asked the
maximum current we need which is linked to the efficiency machine: 5micro
for proton and 6 micro for neutron)
- Need one guy in charge of the luminosity monitor during the experiment


Gagik has presented some results about the calibration of the left
spectrometer (other elog report), Alexander about the DAQ (in attachment)
and Astrid and me about the LED studies (other elog report). 

Catherine
Attachment 1: alex.ppt  56 kB
ELOG V3.1.4-unknown